OpenStudy (anonymous):

Find 3 consecutive odd integers whose sum is 225, explain please

OpenStudy (anonymous):

sorry i ment 255

OpenStudy (zarkon):

\[n+n+2+n+4=225\] \[3n+6=225\] \[3n=219\] \[n=73\] so 73,75,77

OpenStudy (karatechopper):

x=1st number x+2=2nd number x+4=3rd number x+x+2+x+4=225 combine like terms 3x+6=225 3x-6+6=225-6 3x=219 3x/3=219/3 x=73:first number x+2=75:2nd number x+4:3rd number

OpenStudy (karatechopper):

now if tht is 255 then.....

OpenStudy (zarkon):

n+n+2+n+4=255 3n+6=255 3n=249 n=83 so 83,85,87

OpenStudy (karatechopper):

x=1st x+2=2nd x+4=3rd x+x+2+x+4=255 combine like terms 3x+6=255 3x+6-6=255-6 3x=249 3x/3=249/3 x=83:1st x+2=85:2nd x+4=87:3rd

OpenStudy (anonymous):

why did you put x+2 and x + 4?

OpenStudy (karatechopper):

well i used to wonder the same thing, but my teacher just said tht when u have to add something odd together count by twos and u will get the right answer..

OpenStudy (anonymous):

ohh i get it, so if it was 4 numbers the 4th one will be x+6 right?

OpenStudy (karatechopper):

yes sir!

OpenStudy (anonymous):

thanks! :)